# Optimal probe states for single-mode quantum target detection in arbitrary object reflectivity

### v1

@@ IV Summary
In this section, we consolidate the study and results from the preceding sections. Expanding on the preceding discussion, we can distill it into two scenarios. In the case where $\bar{n}\geq\bar{n}_{env}$, the characteristics of $\hat{\rho}_{pr}$ wield a dominant influence on the received states throughout the entire range of $r$. Consequently, OPSs consistently shift towards PNSS as $r$ increases. Conversely, when $\bar{n}<\bar{n}_{env}$, OPSs display features that allow us to classify them into two distinct regions. For $r<r_{T}$, the coherence of OPSs surpasses that of coherent states, presenting phase-squeezed states that predominantly utilize the off-diagonal term of $\hat{\rho}_{pr}$ as the information source for executing QTD. When $r>r_{T}$, $\hat{\rho}_{pr}$ recapture the domination of $\hat{\rho}_{recv}$, the OPSs turn back to PNSSs.
At the specific point $r=r_{T}$, identified as a transition point, OPSs transform from phase-squeezed states to PNSSs; in other words, there is a demarcation point for the main information in QTD at $r=r_{T}$. Intuitively, as the ratio of $\bar{n}_{env}/\bar{n}$ increases with the growth of $r_{T}$, it implies that $\hat{\rho}_{recv}$ requires a greater composition of probe states to counteract the influence of $\hat{\rho}_{env}$. This speculation is also illustrated in Fig.5. Interestingly, OPSs assume the form of coherent states at $r=r_{T}$, resulting in the absence of quantum advantage at this specific point. This occurrence presents a potential strategy for the target or jammer to weaken the impact of QTD.
@@ V Conclusion
In conclusion, this study utilizes an optimization algorithm to pinpoint OPSs throughout the entire reflectivity region. Furthermore, by conducting a comparative analysis with alternative methods employed for OPSs assessment, we offer a comprehensive physical interpretation of the observed phenomena. In the future, the analytical approach could be expanded to encompass two- or even higher-mode QTD, aiming to identify OPSs with entanglement and contribute to a more comprehensive theoretical model of quantum illumination and quantum target detection.

### v2

@@ IV Summary
In this section, we consolidate the study and results from the preceding sections. Based on the preceding discussion, we can distill OPSs into two scenarios. In the case where $\bar{n}\geq\bar{n}_{env}$, the characteristics of $\hat{\rho}_{pr}$ wield a dominant influence on the received states throughout the entire range of $r$. Consequently, OPSs consistently shift towards PNSS as $r$ increases. Conversely, when $\bar{n}<\bar{n}_{env}$, OPSs display features that allow us to classify them into two distinct regions. For $r<r_{T}$, the coherence of OPSs surpasses that of coherent states, presenting phase-squeezed states that predominantly utilize the off-diagonal term of $\hat{\rho}_{pr}$ as the information source for executing QTD. When $r>r_{T}$, $\hat{\rho}_{pr}$ recapture the domination of $\hat{\rho}_{recv}$, the OPSs turn back to PNSSs. In addition, regardless of whether $\bar{n}\geq\bar{n}_{env}$ or $\bar{n}<\bar{n}_{env}$, the OPSs are near-coherent states and PNSSs (Eq.12) as $r$ tends towards 0 and 1, respectively.
At the specific point $r=r_{T}$, identified as a transition point, OPSs transition from phase-squeezed states to PNSSs. In other words, $r=r_{T}$ marks a demarcation point for the primary information utilized in QTD. When $r<r_{T}$, the off-diagonal term of OPSs takes main effect in distinguishing between the $H_{0}$ and $H_{1}$. For $r>r_{T}$, it transitions to the diagonal term of OPSs.
Intuitively, as the ratio of $\bar{n}_{env}/\bar{n}$ increases with the growth of $r_{T}$, it implies that $\hat{\rho}_{recv}$ requires a greater composition of probe states to counteract the influence of $\hat{\rho}_{env}$. This speculation is also illustrated in Fig.5. Interestingly, OPSs assume the form of coherent states at $r=r_{T}$, resulting in the absence of quantum advantage at this specific point. This occurrence presents a potential strategy for the target or jammer to weaken the impact of QTD.
@@ V Conclusion
In conclusion, this study utilizes an optimization algorithm to identify single-mode OPSs across the entire range of target reflectivity, thereby complementing the existing knowledge of OPSs in non-extreme detection conditions. Additionally, through a comparative analysis with alternative methods for OPSs assessment, we provide a clear physical interpretation of the observed phenomena. For future work, expanding the analytical approach to include two- or even higher-mode QTD could facilitate the identification of OPSs with entanglement and contribute to a more comprehensive theoretical model of quantum illumination and quantum target detection. On the other hand, from the perspective of the target and jammer, their objective is to diminish the performance of the prober’s quantum advantage. In this scenario, the jammer could emit various possible noise quantum states arbitrarily to reduce the information obtained by the prober. Therefore, from the perspective of both the target and the jammer, identifying an optimal jammer (noise) quantum state will also be a crucial future task in defending against quantum target detection.
@@ Acknowledgements
Appendix
In this section, we introduce the optimal algorithm, sequential quadratic programming (SQP), used to evaluate the OPSs in this study. SQP is a numerical optimization algorithm utilized to minimize a nonlinear objective function while adhering to equality and inequality constraints.
Fig.6 presents the flow chart of SQP. Consider a parent problem that needs to find the solution $x_{k}$ to minimize a target function $f(x)$. Newton’s method provides an iterative approach for finding this solution. However, when the problem includes constraints such as $h(x)\geq 0$ and $g(x)=0$, SQP suggests considering a quadratic programming (QP) subproblem to calculate the Newton step direction $dx_{k}$ in order to generate a better approximation $x_{k+1}$. In QP subproblem, the target function is transformed to
$f(x_{k})+\nabla f(x_{k})^{T}dx_{k}+\frac{1}{2}dx_{k}^{T}\nabla_{xx}^{2}\mathcal{L}(x_{k},\lambda_{k},\sigma_{k})dx_{k},$ (17)
with the constraints that be considered to the first-order term of Taylor expansion
$h(x_{k})+\nabla h(x_{k})^{T}dx_{k}$ $\geq 0,$ (18)
$g(x_{k})+\nabla g(x_{k})^{T}dx_{k}$ $=0.$
We begin the algorithm by initializing a set of parameters, denoted as $[x_{0},\lambda_{0},\sigma_{0}]$. These initial values are then fed into the QP subproblem to compute the first Newton step, $dx_{0}$, using the Lagrange multiplier method. Subsequently, the parameter set in the parent problem is updated as $[x_{1},\lambda_{1},\sigma_{1}]^{T}=[x_{0},\lambda_{0},\sigma_{0}]^{T}+dx_{0}$. This updated parameter set serves as the initial condition for the second round of the QP subproblem and is utilized to compute the subsequent Newton step, $dx_{1}$. This process is repeated iteratively, updating the parameter set as $[x_{k+1},\lambda_{k+1},\sigma_{k+1}]^{T}=[x_{k},\lambda_{k},\sigma_{k}]^{T}+dx_{k}$, until the parent problem converges. Finally, the resulting $x_{k}$ values that minimize $f(x)$ are obtained, providing the solution to the optimization problem.
In our case, the target function $f(x)$ is the error probability (Eq.1), while the constraint $g(x)$ is defined by the normalization condition, Eq.6, and the finite mean photon number of probe states, Eq.7. Due to computational limitations, the dimension of the calculated density matrix is restricted to 8, spanning from the vacuum state $|0\rangle$ to the number state $|7\rangle$. By evaluating the algorithm with these equations and conditions, we can determine the optimal probe states under arbitrary detection conditions.